OpenStudy (anonymous):

Find the area of the region that lies inside the first curve and outside the second curve

OpenStudy (anonymous):

\[r= 12 \sin \theta , r=6\]

OpenStudy (anonymous):

i know that \[A= \frac{ 1 }{ 2 }\int\limits_{a}^{b}(r:1st)^2-(r:2nd)^2 dtheta\] but im not sure how to find a and b on the integral

OpenStudy (anonymous):

Do i set the 2 equal to each other and solve for theta to find a and b?

OpenStudy (perl):



OpenStudy (perl):

thats the area you want?

OpenStudy (perl):

first solve $$\Large { r= 12 \sin \theta , r=6\\ 6 = 12 \sin \theta \\ \frac 1 2 = \sin \theta \\ \theta = \pi /6 ~, ~5 \pi / 6 } $$

OpenStudy (anonymous):

it wouldn be just neg and pos pi/6

OpenStudy (anonymous):

idk, this is really hard!

OpenStudy (perl):

$$ \Large { \frac12\int_{\pi/6}^{5\pi/6}\frac12 [(12 \sin \theta)^2 - 6^2]~ d\theta } $$

OpenStudy (perl):

sin theta = 1/2 , the solution comes from the unit circle .

OpenStudy (anonymous):

ok

OpenStudy (perl):